# SynPO: Synergizing Descriptiveness and Preference Optimization for Video Detailed Captioning

### v1

@@ Appendix C Implementation Details for Video Captioning
  • Video Detailed Captioning (VDC) [chai2024auroracap] transforms the matching between two paragraphs into a set of question-answer pairings. It first generates some question-answer pairs based on the ground truth captions, then derive corresponding answers one by one from the generated captions, and finally perform matching. The process is automatically evaluated with the LLM involvement in each step.
  • Video Detailed Description (VDD) [li2023videochat] is a multimodal benchmark designed to evaluate models’ ability to generate temporally coherent, semantically rich, and contextually precise natural language descriptions of video content, integrating visual, and textual modalities through datasets with fine-grained captions to challenge cross-modal reasoning, dynamic scene understanding, and long-term temporal modeling in video-language tasks. Notably, it utilizes LLM to score the similarity between ground-truth caption and generated caption.
  • Microsoft Research Video to Text (MSRVTT) [xu2016msr] is a large-scale multimodal benchmark designed to evaluate models’ ability to generate temporally coherent and contextually rich textual descriptions of video content, comprising 10,000 video clips annotated with 20 English sentences each via crowdsourcing, and featuring standardized train/validation/test splits across 20 diverse categories to challenge cross-modal reasoning and dynamic scene understanding in video-language tasks.
  • VATEX [wang2020vatexlargescalehighqualitymultilingual] is a large-scale multilingual multimodal benchmark designed for video captioning and cross-lingual machine translation tasks, comprising 41,250 video clips annotated with 825,000 English-Chinese subtitles (206,000 aligned pairs), emphasizing cross-modal reasoning, temporal coherence, and linguistic diversity to evaluate models’ ability to generate context-aware descriptions and leverage visual-spatial cues for accurate multilingual translation.

### v2

@@ Appendix C Implementation Details for Video Captioning
  • Video Detailed Captioning (VDC) [undefj] transforms the matching between two paragraphs into a set of question-answer pairings. It first generates some question-answer pairs based on the ground truth captions, then derive corresponding answers one by one from the generated captions, and finally perform matching. The process is automatically evaluated with the LLM involvement in each step.
  • Video Detailed Description (VDD) [undefak] is a multimodal benchmark designed to evaluate models’ ability to generate temporally coherent, semantically rich, and contextually precise natural language descriptions of video content, integrating visual, and textual modalities through datasets with fine-grained captions to challenge cross-modal reasoning, dynamic scene understanding, and long-term temporal modeling in video-language tasks. Notably, it utilizes LLM to score the similarity between ground-truth caption and generated caption.
  • Microsoft Research Video to Text (MSRVTT) [undefaay] is a large-scale multimodal benchmark designed to evaluate models’ ability to generate temporally coherent and contextually rich textual descriptions of video content, comprising 10,000 video clips annotated with 20 English sentences each via crowdsourcing, and featuring standardized train/validation/test splits across 20 diverse categories to challenge cross-modal reasoning and dynamic scene understanding in video-language tasks.
  • VATEX [undefaaq] is a large-scale multilingual multimodal benchmark designed for video captioning and cross-lingual machine translation tasks, comprising 41,250 video clips annotated with 825,000 English-Chinese subtitles (206,000 aligned pairs), emphasizing cross-modal reasoning, temporal coherence, and linguistic diversity to evaluate models’ ability to generate context-aware descriptions and leverage visual-spatial cues for accurate multilingual translation.